OpenStudy (anonymous):

Caleb and Sergio stacked boxes on a shelf. Caleb lifted 14 boxes and Sergio lifted 12 boxes. The boxes that Sergio lifted each weighed 10 lb more than the boxes Caleb lifted. Which expression represents the total number of pounds Sergio lifted? Let c represent the weight of each box Caleb lifted. A. 120c B. c + 10 C. 12(c + 10) D. 10(c − 12)

OpenStudy (ericanoel912):

12(c+10) :)

OpenStudy (anonymous):

Thanks,i got 4 more if you want to figure them out.

OpenStudy (ericanoel912):

np and sure :)

OpenStudy (ericanoel912):

you can PM me them if that's easier

OpenStudy (anonymous):

Thank you so much. Two stores sell the same television for the same original price. Store A advertises that the television is on sale for 30 percent off the original price. Store B advertises that it is reducing the television’s price by 250 dollars. When Allison compares the sale prices of the television in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the television’s original price. A. 0.3p=p-250 B. 0.7p=p-250 C. 0.3p=p+250 D. 0.7(p-250)=p

OpenStudy (anonymous):

those question marks are just commas.ill pm you,if you cant figure them out now.

OpenStudy (ericanoel912):

It's A the 1st one is 30% of original so it's .3 x p the 2nd one is $250 less than original, so it's p-250

OpenStudy (anonymous):

These 3 questions left might be a little easier. Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel? MPH:

OpenStudy (ericanoel912):

I can't get this 1 sry, next one?

OpenStudy (anonymous):

Gretchen and Ezia received equal scores on a test made up of multiple choice questions and an essay. Gretchen got 18 multiple choice questions correct and received 19 points for her essay. Ezia got 15 multiple choice questions correct and received 31 points for her essay. How many points was each multiple choice question worth? Points:

OpenStudy (ericanoel912):

srry can't get this either!! Maybe submit to mathematics forum

OpenStudy (anonymous):

k,thank you bro

OpenStudy (ericanoel912):

lol EricA